OpenStudy (anonymous):

Cashews cost $4 per pound and peanuts cost $1 per pound. How many pounds of peanuts should be used with the cashews to get a mixture of 24 pounds that will cost $2 per pound?

OpenStudy (anonymous):

I know you are going to need a system of equations, I will attempt to set it up for you let y=cashews let x=penauts x+y=2 x+4y=(24)(2) i believe that will work out for you and if not i am going to have to find some notes

OpenStudy (anonymous):

change the top one to x+y=24

OpenStudy (anonymous):

yea if you change the first one its right

OpenStudy (anonymous):

That one was reaching way back!